OpenStudy (anonymous):

which points are the best approximations of the relative maximum and minimum of the function f(x)=x^3+3x^2-9x-8

OpenStudy (anonymous):

theser are the answer options HELP!!

OpenStudy (amoodarya):

find f' find roots of f'=0 check them by f' sign table or ,sign of f'' to find out max or min

OpenStudy (anonymous):

find f'(x) and then equate to 0 to find the critical points

OpenStudy (anonymous):

then check f"(x) at each point.

OpenStudy (anonymous):

would it be B??

OpenStudy (anonymous):

i really need please pls someone

OpenStudy (anonymous):

we have not solved yet. you solve t and show ,we will guide you.

OpenStudy (anonymous):

i just dont even know where to start..

OpenStudy (anonymous):

differentiate with respect to x

OpenStudy (anonymous):

\[f \prime \left( x \right)=3x^2+6x+9\]equate it to zero and find values of x

OpenStudy (anonymous):

correction it is -9 not +9

OpenStudy (anonymous):

so f(x) = 9 when equated to 0??

OpenStudy (anonymous):

no \[f \prime \left( x \right)=0~means~3x^2+6x-9=0\] divide by 3

OpenStudy (anonymous):

oh so x^2 + 2x - 3 = 0

OpenStudy (anonymous):

make factors

OpenStudy (anonymous):

x^2+(3-1)x-3=0 x^2+3x-1x-3=0

OpenStudy (anonymous):

how do u make factors

OpenStudy (anonymous):

3-1=-3 3*-1=-3

OpenStudy (anonymous):

correction 3-1=2

OpenStudy (anonymous):

coefficient of x^2*constant term is same as the product of two terms (made from coefficient of x)

OpenStudy (anonymous):

ok i still dont understand ugh

OpenStudy (anonymous):

coefficient of x^2=1 constant term=-3 product=1*-=-3 if the product is negative always subtract something to make the middle term here 2=3-1 product of 3*-1=-3

OpenStudy (anonymous):

i am leaving for sometime . i will be back to help you after 5 minutes.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

nevermind i figured it out

OpenStudy (anonymous):

i am here.

OpenStudy (anonymous):

have you found f"(x)

OpenStudy (anonymous):

Have you calculated critical points? \[f \prime \prime \left( x \right)=6x+6\]

OpenStudy (anonymous):

x(x+3)-1(x+3)=0 (x+3)(x-1)=0 x+3=0,x=-3 or x-1=0,x=1

OpenStudy (anonymous):

\[f \prime \prime \left( -3 \right)=?\]